# TacoGFN: Target-conditioned GFlowNet for Structure-based Drug Design

### v5

@@ Abstract
Searching the vast chemical space for drug-like and synthesizable molecules with high binding affinity to a protein pocket is a challenging task in drug discovery. Recently, molecular deep generative models have been introduced which promise to be more efficient than exhaustive virtual screening, by directly generating molecules based on the protein structure. However, since they learn the distribution of a limited protein-ligand complex dataset, the existing methods struggle with generating novel molecules with significant property improvements. In this paper, we frame the generation task as a Reinforcement Learning task, where the goal is to search the wider chemical space for molecules with desirable properties as opposed to fitting a training data distribution. More specifically, we propose TacoGFN, a Generative Flow Network conditioned on protein pocket structure, using binding affinity, drug-likeliness and synthesizability measures as our reward. Empirically, our method outperforms state-of-art methods on the CrossDocked2020 benchmark for every molecular property (Vina score, QED, SA), while improving the generation time by multiple orders of magnitude. TacoGFN achieves $-8.82$ in median docking score and $52.63\%$ in Novel Hit Rate.
Machine Learning, ICML
Figure 1: Example of molecules generated for target 5w2g by DecompDiff (Guan et al., 2023b) - a current SOTA baseline (a1) and our model (b1), presented alongside the with the most similar molecule from the training set (a2, b2). Relevant properties of generated molecules are listed: $\uparrow$ denotes higher is better, $\downarrow$ denotes lower is better.
@@
1 Introduction
Structure-based drug design (SBDD) leverages target protein structures to search for high affinity drug molecules. Due to the growing availability of protein structures from ML protein structure prediction methods (Jumper et al., 2021), and many novel targets identified from high-throughput perturbation experiments (Replogle et al., 2022), SBDD is becoming an increasingly powerful approach in drug discovery. It currently takes 13-15 years and between US \$2 billion and \$3 billion for a single drug to be developed and approved (Pushpakom et al., 2018). The substantial expense and time of drug development not only impose a significant burden on healthcare systems but also amplify societal risks during global health crises, such as the COVID-19. There is an urgent need to expedite the design of novel drug candidates for new protein targets.
Traditionally, molecular docking has been used to virtually screen libraries of molecules for interaction with a target protein. Its efficacy is impeded by the exhaustive nature of its search, and by the high computational cost of molecular docking. To overcome this challenge, recently deep generative models (Luo et al., 2021; Peng et al., 2022; Guan et al., 2023b) have been proposed to design molecules (ligands) conditioned on a pocket structure. These methods learn the distribution of a training dataset of protein-ligand complex structures.
However, due to the high cost of the experiments, the size of the training datasets for SBDD, i.e. high-quality protein-ligand binding structure data, is relatively small. PDBBind (Liu et al., 2014), the dataset on which the standard CrossDocked2020 benchmark (Francoeur et al., 2020) is based on, contains only 19,443 protein-ligand complexes. After removing common biomolecules and duplicates, only 4,200 unique drug-like molecules remain (Powers et al., 2023). This is only a tiny fraction of the entire chemical space of drug-like molecules that is assumed to consist of $\sim 10^{60}$ molecules (Lipinski et al., 1997). As a consequence, existing methods of SBDD based on distribution learning have struggled to generate novel molecules with significantly improved properties (Lee et al., 2023), since they generate molecules with high structural similarity to the training set (Guan et al., 2023b).
In this paper, we frame the task of SBDD as a Reinforcement Learning (RL) task, where the goal is to search the chemical space for molecules with desirable molecular properties as opposed to fitting a training data distribution. We propose TacoGFN, a Target Conditioned Generative Flow Network that generates molecules conditioned on a given protein pocket structure, guided by rewards based on predicted affinity of the molecule to the pocket, drug-like properties and synthesizability of the molecule. We introduce a docking score predictor which leverages pre-trained pharmacophore representation to efficiently evaluate millions of diverse protein-ligand pairs during training. We performed an experimental evaluation on the standard CrossDocked benchmark dataset and demonstrated that TacoGFN clearly outperforms state-of-the-art methods. Our method achieved up to $-8.82$ in median docking score while obtaining far superior values of drug-likeness and synthesizability. Furthermore, TacoGFN reaches a Novel Hit Rate of $52.63\%$, which is a dramatic improvement over the $7.94\%$ obtained by the best performing baseline model.
To summarize, the main contributions of this paper are:
  • We point out the limitations of previous distribution learning approaches and propose to frame pocket-conditioned molecule generation as a multi-objective RL task.
  • We propose TacoGFN, which conditions a GFlowNet for learning a family of molecule distributions conditioned on protein pocket structures.
  • We develop a docking score predictor with improved generalization capabilities using a pre-trained pharmacophore representation, enabling fast affinity evaluation.
  • TacoGFN clearly outperforms the current SOTA on the CrossDocked2020 benchmark and demonstrates the benefit of our approach over existing distribution learning based methods.
@@ 2 Related Work
Structure-based drug design aims to sample drug-like molecules that fit a given target protein pocket. LiGAN (Masuda et al., 2020) uses 3D CNNs to encode the protein pocket structure and predict atom densities from the encoded latent space. 3DSBDD (Luo et al., 2021) and Pocket2Mol (Peng et al., 2022) adopt an auto-regressive approach to generate molecules atom by atom. Other methods such as FLAG (Zhang et al., 2023b) and DrugGPS (Zhang & Liu, 2023) build molecules fragment by fragment to leverage the chemical prior. A very recent line of research employs diffusion models (Guan et al., 2023a, b; Schneuing et al., 2023) for SBDD. TargetDiff (Guan et al., 2023a) is a diffusion-based method which generates atom coordinates and atom types in a non-autoregressive way, and bonds are generated in a post-processing step. DecompDiff (Guan et al., 2023b) is a diffusion model which generates both atoms and bonds with decomposed priors, which reflect the natural decomposition of a ligand molecule into arms and scaffold. DecompDiff has achieved state-of-the-art performance on the CrossDocked2020 benchmark. A recent analysis paper (Harris et al., 2023) questions the assumption that explicit 3D modelling of the ligand improves performance, after finding a much higher occurrence of physical violations and fewer key interactions in molecules generated using 3D modelling. In our work, we generate molecules in 2D space to vastly reduce the search space and compute time; Our approach is shown to effectively leverage the target pocket structure for generating molecules with high affinity.
@@
Goal-directed molecule generation aims to generate molecules that satisfy certain optimization goals such as strong binding affinity and high drug-likeliness. Reinforcement Learning (RL) methods such as ReLeaSE (Olivecrona et al., 2017), MolDQN (Zhou et al., 2019) and REINVENT (Blaschke et al., 2020) have been proposed to guide the generation of molecules toward desirable properties. MORLD (Jeon & Kim, 2020) and MoleGuLAR (Goel et al., 2021) combine RL and docking calculations to design novel ligands. MOOD (Lee et al., 2023) incorporates out-of-distribution and property-guided exploration in diffusion models for goal-directed molecule generation. Existing goal-directed methods are constrained by a fixed objective function. They are trained to generate high-affinity molecules for a single target pocket, and therefore, a new model is needed for each different protein pocket. In contrast, TacoGFN is conditioned on a protein pocket graph and is trained to generate high-affinity molecules for any pocket structure.
@@ 3 GFlowNet Preliminaries
Generative Flow Networks (GFlowNets, GFN) (Bengio et al., 2021) learn a stochastic policy $\pi$ for generating a combinatorial object (such as molecular graph) $x\in\mathcal{X}$. The probability of constructing $x$, denoted as $\pi(x)$, is trained to be proportional to a non-negative reward function $R:\mathcal{X}\mapsto\mathbb{R}^{+}$ defined on the space $\mathcal{X}$. This property of GFlowNet is ideal for generating diverse molecules with desirable properties. Conditional GFlowNet introduced in (Jain et al., 2023) simultaneously model a family of reward functions. Each conditional information, denoted as $c\in C$, induces a unique reward function $R(x|c)$. In our work, we adopt conditional GFlowNet for SBDD settings, by encoding target pocket structure as condition $c$. Thus, a single GFlowNet models high-reward molecule distribution across all protein pockets.
@@
The forward policy $P_{F}(s^{\prime}|s,c)$ of a GFlowNet represents the probability distribution of reaching state $s^{\prime}$ from state $s$ conditioned on context $c$. Partition function $Z(c)$ is the sum of the rewards $R(x|c)$ for all objects $x\in\mathcal{X}$. We adopt the Trajectory Balance objective from Equation  1 to efficiently learn a forward policy $P_{F}$ that generates object $x$ with probability proportional to its reward $R(x|c)$ (Malkin et al., 2023).
$\mathcal{L}_{TB}(\tau,c;\theta)=\left(\log\frac{Z_{\theta}(c)\prod_{s{\rightarrow}s^{\prime}\in\tau}P_{F_{\theta}}(s^{\prime}|s,c)}{R(x|c)}\right)^{2},$ (1)
$\theta$ represents the learnable parameters. Our goal is to learn a policy $\pi(x|c)\approx\frac{R(x|c)}{Z(c)}$ across all molecules $x\in\mathcal{X}$ and protein pocket contexts $c\in\mathcal{C}$.
@@ 4 TacoGFN
Problem definition. The goal of structure-based drug design is to generate molecules with both desirable properties and strong binding affinity with respect to a given protein pocket structure. The pocket structure is denoted as $p$ and will be represented as a $K$-nearest neighbor (KNN) residue graph. The ligand is denoted as $l$ and will be represented either as an atom graph or a fragment graph. We define a reward function $R(l|p)$ based on a molecule’s predicted docking score, drug-likeliness and synthesizability. Our goal is to learn a molecule generation policy that constructs molecules $l$ given protein structure $p$ with probability matching their reward, such that $\pi(l|p)\propto R(l|p)$.
Since TacoGFN’s training objective is based on the reward and not a fixed dataset, it explores a much greater chemical space and discovers novel molecules beyond the known data distribution. In addition, by matching the generation probability to the reward, TacoGFN can discover molecules that improve multiple objectives compared to the training data distribution. In summary, TacoGFN explores the wider chemical space for high reward candidates instead of memorizing a small set of known binding molecules.
Method overview. TacoGFN is a single conditioned generative model that generalizes over protein pockets. It explores the greater chemical space to generate high-affinity molecules with properties desirable as a drug candidate. The protein pocket is first featurized as a $K$-nearest neighbor residue graph and encoded using GVP-GNN (Jing et al., 2021) (section: 4.1). Then, we use this pocket embedding as the condition for GFlowNet molecule generation (section 4.2). Finally, we reward generated molecules using our docking score prediction model which leverages protein-ligand interaction priors (section 4.3). The high-level architecture of our SBDD framework is illustrated in Figure 2.
@@ 4.2 Pocket conditioned GFlowNet
Molecule representation. During molecular generation, we represent ligands as a 2D molecular graph $\mathcal{G}^{L}=(\mathcal{V}^{L},\mathcal{E}^{L})$ with node $v^{L}_{i}\in\mathcal{V}^{L}$ representing a molecule fragment, and directional edges $e^{L}_{ij}\in\mathcal{E}^{L}$ indicating the attachment atom of fragment $v^{L}_{i}$ that connects to fragment $v^{L}_{j}$. Since a molecule’s desirability as a drug in the real world is independent of its predicted 3D conformation, the 2D representation of a ligand here is appropriate.
Fragment vocabulary construction. TacoGFN generates molecules by adding one molecular fragment at a time. To create the vocabulary of fragments used, we extract common fragments from a chemical database in a data-driven and chemically valid way. To obtain a fragment vocabulary, we first apply BRICS decomposition (Degen et al., 2008) to 250k ZINC20 (Irwin et al., 2020) molecules. BRICS breaks molecules via retrosynthetic rules and provides synthetically accessible building blocks, i.e. molecules that are easy to prepare (Seo et al., 2023). To reduce the fragment vocabulary size, we further break all single bonds connecting a heavy atom to a ring structure. Next, we retain the fragments that occur in more than 50 (0.02%) of the molecules in our ZINC sample. Finally, we merge identical fragment graphs. Attachment points are atoms with a broken bond resulting from the molecule decomposition. This results in a set of 475 building block-like fragments. As our fragments are mined from a synthetically accessible virtual library, TacoGFN can generate more synthesizable and chemically valid molecules compared to atom-based construction methods.
Molecular generation framework. We formulate molecular generation as a sequential decision process and implement it using a GFlowNet. At the $t$-th step, the policy $\pi$ samples an action $a$ depending on the current molecule state $\mathcal{G}^{L}_{t}$ and pocket embedding $\mathbf{h}_{\mathcal{G}^{\mathcal{P}}}$. The transition function $\mathcal{T}$ is a deterministic function which applies the action to the molecular graph at step $t$ to produce a molecule graph at step $t+1$.
$a_{t}$ $\sim\pi(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ (2)
@@
Previous autoregressive models often formulate action prediction as a supervised task, where the goal is to predict the correct ground truth actions obtained from masked molecules (Peng et al., 2022). Instead, our molecule generation policy $\pi$ aims to generate molecules with probability $P(\mathcal{G}^{L}|\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ proportional to the reward.
@@
We use a graph transformer (Yun et al., 2020) to model our policy $\pi(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$, by taking the partially constructed molecular graph at the $t$-th time step $\mathcal{G}^{L}_{t}$ and the pocket embedding $\mathbf{h}_{\mathcal{G}^{\mathcal{P}}}$ as input. The input feature $\mathbf{h}^{L(0)}_{i}$ of molecular node $v^{L}_{i}$ is a one-hot encoding of the node’s fragment type. The molecular edge input feature $\mathbf{e}_{ij}^{L(0)}$ is a one-hot mapping of the attachment atom index in node $v^{L}_{i}$ which connects to node $v^{L}_{j}$. We add an additional virtual conditioning node $\mathbf{h}^{V(0)}$, featurized using pocket embedding $\mathbf{h}_{\mathcal{G}^{\mathcal{P}}}$, to the graph $\mathcal{G}^{L}_{t}$ (Pham et al., 2017). This virtual node is connected to all other nodes and serves as a graph-level node to provide pocket information. After $N$ Transformer layers, the set of final node embeddings $\{\mathbf{h}^{L(N)}_{i}\}$ and edge embeddings $\{\mathbf{e}^{L(N)}_{ij}\}$ are obtained. The final graph level embedding $\mathbf{g}^{L(N)}$ is obtained via the concatenation of the global average pooling of the node embeddings and the final virtual node embedding $\mathbf{h}^{V(N)}$, as seen in Equation 4.
Using these final molecular graph embeddings, we sample from the following three action types defined in previous works on fragment-based molecular generation (Bengio et al., 2021; Jin et al., 2018): 1) Fragment addition is a node action that connects the current node to a new node via a new edge. An MLP is applied on the node embeddings $\mathbf{h}^{L(N)}_{i}$ to produce logits over the fragment vocabulary for every node in the molecular graph. 2) Attachment specification is an edge action that specifies the attachment atom in a node (fragment) which will form a single bond with the other fragment. An MLP is applied on each edge embedding $\mathbf{e}^{L(N)}_{ij}$ to produce logits over attachment point choices. 3) Stop construction is a graph action that marks the finish of a molecule. The logit is produced from a single MLP output based on the final graph embedding $\mathbf{g}^{L(N)}$. All logits are concatenated and scaled into probabilities using the softmax function, and an action is sampled from the distribution (see Equations in B.1). The same process is repeated for each time step until the stop construction action is sampled.
Reward function. A drug candidate must not only have a high affinity to the pocket but also satisfy drug-like properties and synthetizability requirements to be selected for experimental validation. We design a reward function which combines all three aspects, using QED (Bickerton et al., 2012) as a measure of drug-likeliness, SA (Ertl & Schuffenhauer, 2009) as a measure of ease-of-synthesizability, and predicted docking score as a measure of affinity between ligand and protein (see details in Appendix B.2). Computing the reward is very fast, for example, computing it for a batch of 64 protein-ligand pairs took under 0.15 seconds.

### v6

# Target-conditioned GFlowNet for Structure-based Drug Design

@@ Abstract
Searching the vast chemical space for drug-like molecules that bind with a protein pocket is a challenging task in drug discovery. Recently, structure-based generative models have been introduced which promise to be more efficient by learning to generate molecules for any given protein structure. However, since they learn the distribution of a limited protein-ligand complex dataset, structure-based methods do not yet outperform optimization-based methods that generate binding molecules for just one pocket. To overcome limitations on data while leveraging learning across protein targets, we choose to model the reward distribution conditioned on pocket structure, instead of the training data distribution. We design TacoGFN, a novel GFlowNet-based approach for structure-based drug design, which can generate molecules conditioned on any protein pocket structure with probabilities proportional to its affinity and property rewards. In the generative setting for CrossDocked2020 benchmark, TacoGFN attains a state-of-the-art success rate of $56.0\%$ and $-8.44$ kcal/mol in median Vina Dock score while improving the generation time by multiple orders of magnitude. Fine-tuning TacoGFN further improves the median Vina Dock score to $-10.93$ kcal/mol and the success rate to $88.8\%$, outperforming all optimization-based methods.
1 Introductions
Structure-based drug design (SBDD) leverages target protein structures to search for high-affinity drug molecules. Due to the growing availability of protein structures from ML protein structure prediction methods (Jumper et al., 2021), and many novel targets identified from high-throughput perturbation experiments (Replogle et al., 2022), SBDD is becoming an increasingly powerful approach in drug discovery. It currently takes 13-15 years and between US \$2 billion and \$3 billion for a single drug to be developed and approved (Pushpakom et al., 2018). The substantial expense and time of drug development not only impose a significant burden on healthcare systems but also amplify societal risks during global health crises, such as COVID-19. There is an urgent need to expedite the design of novel drug candidates for new protein targets.
Traditionally, virtual screening has been used to discover binding molecules by predicting supramolecular interactions between ligands and a target protein with molecular docking. Its efficacy is impeded by the exhaustive nature of its search, and by the high computational cost of molecular docking. To overcome this challenge, many optimization-based methods (Bengio et al., 2021; Fu et al., 2022; Lee et al., 2023; Reidenbach, 2024) have been proposed to generate high-affinity molecules for one protein pocket only. These methods typically do not take protein structure as input and use docking measures as the reward. Recently, structure-based generative models (Luo et al., 2021; Peng et al., 2022; Guan et al., 2023b) have been proposed to design molecules (ligands) conditioned on the pocket structure. These methods learn the distribution of a training dataset of protein-ligand complexes. Structure-based generative model has the advantage of learning generalized protein-ligand interaction patterns by leveraging different pocket structures during training.
However, due to the high cost of the experiments, the size of the training datasets for SBDD, i.e. high-quality protein-ligand binding structure data, is relatively small. PDBBind (Liu et al., 2014), the underlying dataset of the standard CrossDocked2020 benchmark (Francoeur et al., 2020), contains only 19,443 protein-ligand complexes. After removing common biomolecules (lipids, peptides, carbohydrates, and nucleotides) and duplicates, only 4,200 unique drug-like molecules remain (Powers et al., 2023). This is only a tiny fraction of the entire chemical space of drug-like molecules that is assumed to consist of $\sim 10^{60}$ molecules (Lipinski et al., 1997). As a consequence, existing structure-based generative models relying on data-distribution learning model only a very small part of the overall chemical space and have struggled to generate novel molecules with significantly improved properties (Lee et al., 2023); Therefore, existing structure-based generative methods do not yet outperform optimization-based methods on a single target.
In summary, many current optimization-based methods cannot leverage learning from across protein pocket structures - they are limited to a reward function based on affinity to one single protein target per model. On the other hand, while structure-based models can learn generalized protein-ligand interaction patterns by training on different pockets, they are restricted to modelling a limited data distribution. Therefore they are unable to effectively explore the wider chemical space for desirable molecules, and under perform optimization-based methods. Our goal in this paper is to both leverage learning across protein target structures, and overcome the challenge of limited data distribution. To the best of our knowledge, TacoGFN is the first RL model to address the challenging task of modelling a family of reward functions induced from all pocket structures.
In this paper, we frame the task of structure-based molecule generation as learning a reward distribution instead of a training data distribution, and adopt GFlowNet Bengio et al. (2021) - an energy-based generative model for generating combinatorial objects. We propose TacoGFN, a Target Conditioned Generative Flow Network that generates molecules conditioned on any given protein pocket structures, guided by affinity to the pocket, drug-likeness and synthesizability measures. This formulation allows us to explore the greater chemical space as we are no longer constrained to a fixed dataset. We performed an experimental evaluation on the standard CrossDocked benchmark dataset and demonstrated that TacoGFN clearly outperforms state-of-the-art structure-based generative methods, improving the success rate to 56.0% from the previous best of 24.5%. TacoGFN shows even stronger performance with fine-tuning (TacoGFN+FT), resulting in -10.93 kcal/mol in median Vina Dock score and up to 88.8% in success rate, outperforming all optimization-based models.
To summarize, the main contributions of this paper are:
  • We point out the limitations of previous data-based distribution learning, and instead propose to frame structure-based molecule generation as learning from the reward distribution. To this end, we propose TacoGFN, the first application of GFlowNet for learning a family of molecule distributions conditioned on protein pocket structures.
  • We introduced a novel pharmacophore-based affinity predictor, with improved generalization capabilities using a pre-trained pharmacophore representation, enabling fast affinity evaluation.
  • TacoGFN clearly outperforms the state-of-the-art structure-based methods on the CrossDocked2020 benchmark and demonstrates the benefit of reward-based distribution learning.
  • TacoGFN+FT outperforms the optimization methods focused on a single pocket only, learning from diverse molecules for various protein pocket structures. TacoGFN+FT attains the best Vina Dock score, high-affinity rate and success rate among all optimization-based methods.
@@ 2 Related Work
Structure-based molecular generation aims to generate high-affinity molecules for any target pocket. They are expected to generalize on previously unseen test pockets, and therefore, do not have access to the docking oracle for the test pockets during inference or training time. The goal for evaluating the quality of molecules generated for unseen protein structures is to measure whether SBDD models have learned generalized protein-ligand interaction patterns during training. Many approaches for this problem setting has been proposed. LiGAN (Masuda et al., 2020) uses 3D CNNs to encode the protein pocket structure and predict atom densities from the encoded latent space. 3DSBDD (Luo et al., 2021) and Pocket2Mol (Peng et al., 2022) adopt an auto-regressive approach to generate molecules atom by atom. Other methods such as FLAG (Zhang et al., 2023b) and DrugGPS (Zhang & Liu, 2023) build molecules fragment by fragment to leverage the chemical prior. A very recent line of research employs diffusion models (Guan et al., 2023a; b; Schneuing et al., 2023) for SBDD. TargetDiff (Guan et al., 2023a) is a diffusion-based method which generates atom coordinates and atom types in a non-autoregressive way, and bonds are generated in a post-processing step. DecompDiff (Guan et al., 2023b) is a diffusion model which generates both atoms and bonds with decomposed priors, which reflect the natural decomposition of a ligand molecule into arms and scaffold. A recent analysis paper (Harris et al., 2023) questions the assumption that explicit 3D modelling of the ligand improves performance, after finding a much higher occurrence of physical violations and fewer key interactions in molecules generated using 3D modelling. In our work, we generate molecules in 2D space to vastly reduce the search space and compute time; Our approach is shown to effectively leverage the target pocket structure for generating molecules with high affinity.
@@
Optimization-based molecular generation aims to generate molecules that satisfy certain optimization goals. Compared to the distribution-based generative model, these methods can be designed to optimize for molecules with strong binding affinity to a particular protein target. Since they have access to a docking oracle during training or inference time, they are not directly comparable to the structure-based generative problem setting. Reinforcement Learning (RL) methods such as ReLeaSE (Olivecrona et al., 2017), MolDQN (Zhou et al., 2019) and REINVENT (Blaschke et al., 2020) have been proposed to guide the generation of molecules toward desirable properties. MORLD (Jeon & Kim, 2020) and MoleGuLAR (Goel et al., 2021b) combine RL and docking calculations to design novel ligands. RGA (Fu et al., 2022) proposes a variant of a genetic algorithm that is guided by reinforcement learning and is pre-trained on multiple protein target structures. MOOD (Lee et al., 2023) incorporates out-of-distribution and property-guided exploration in diffusion models for goal-directed molecule generation. DecompOpt Zhou et al. (2024) combines a pre-trained structure-based equivariant diffusion model with a docking-based greedy iterative optimization loop. EvoSBDD Reidenbach (2024) improves efficiency by performing black-box optimization over the 1D latent space using a docking oracle. In summary, optimization-based molecular generation evaluates the performance of the algorithm on optimizing for seen targets. Compared to the structure-based setting, which requires modelling the molecule distribution conditional to any protein pocket structure and generalizing to unseen protein pockets, optimizing the molecule distribution for a single protein pocket is a less challenging task. We refer readers to a more detailed comparison of the problem setting in Appendix B.
@@ 3 GFlowNet Preliminaries
Generative Flow Networks (GFlowNets, GFN) (Bengio et al., 2021) learn a stochastic policy $\pi$ for generating a combinatorial object (such as molecular graph) $x\in\mathcal{X}$. The probability of constructing $x$, denoted as $\pi(x)$, is trained to be proportional to a non-negative reward function $R:\mathcal{X}\mapsto\mathbb{R}^{+}$ defined on the space $\mathcal{X}$. This property of GFlowNet is ideal for generating diverse molecules with desirable properties. Conditional GFlowNet introduced in (Jain et al., 2023) simultaneously models a family of reward functions. Each conditional information, denoted as $c\in C$, induces a unique reward function $R(x|c)$. In our work, we adopt conditional GFlowNet for SBDD settings, by encoding target pocket structure as condition $c$, where $c$ is a high dimensional representation of the pocket structure. Thus, a single GFlowNet models high-reward molecule distribution across all protein pockets.
@@
Following previous studies, we introduce an exponent $\beta$ to the reward function $R(x|c)$, thus modelling $\pi(x|c,\beta)\propto R(x|c)^{\beta}$. This steers generating probability distribution to focus on the modes of $R(x|c)$, which is crucial for producing candidates that are high in reward. Adjusting $\beta$ allows us to manage the balance between diversity and achieving higher rewards.
The forward transition probability $P_{F}(s^{\prime}|s,c,\beta)$ of a GFlowNet represents the probability distribution of reaching state $s^{\prime}$ from state $s$ conditioned on context $c$ and reward temperature $\beta$. Partition function $Z(c,\beta)$ is the sum of the rewards $R(x|c)^{\beta}$ for all objects $x\in\mathcal{X}$ under context $c$. We adopt the Trajectory Balance objective from Equation  1 to efficiently learn a forward transition policy $P_{F}$ that generates object $x$ with probability proportional to its reward $R(x|c)^{\beta}$ (Malkin et al., 2023).
$\mathcal{L}_{TB}(\tau,c,\beta;\theta)=\left(\log\frac{Z_{\theta}(c,\beta)\prod_{s{\rightarrow}s^{\prime}\in\tau}P_{F_{\theta}}(s^{\prime}|s,c,\beta)}{R(x|c)^{\beta}}\right)^{2},$ (1)
$\theta$ represents the learnable parameters. Our goal is to model the probability $\pi(x|c,\beta)\approx\frac{R(x|c)^{\beta}}{Z(c,\beta)}$ across all molecules $x\in\mathcal{X}$, protein pocket contexts $c\in\mathcal{C}$ and various reward temperatures $\beta$.
@@ 4 TacoGFN
TacoGFN is a structure-based molecular generative model that generalizes over all protein pockets with a single model. By matching the reward distribution instead of limited data distribution, our method explores the greater chemical space to generate high-affinity molecules with properties desirable as a drug candidate. Furthermore, by encoding pocket structure information, TacoGFN and its fine-tuned variant are able to leverage learning from diverse protein pocket structures.
Problem definition. The goal of structure-based drug design (SBDD) problem is to generate molecules with both desirable properties and strong binding affinity with respect to any given protein pocket structure. The goal for SBDD models is learning generalized protein-ligand interaction patterns during training. Therefore, SBDD models take pocket structures as input and are expected to generalize for unseen pocket structures. We refer readers to a detailed contrast of the SBDD problem setting with the optimization setting in Appendix B
The pocket structure is denoted as $P$ and will be represented as a $K$-nearest neighbour (KNN) residue graph. The ligand is denoted as $L$ and will be represented either as an atom graph or a fragment graph. We define a reward function $R(L|P)$ based on a molecule’s predicted docking score, drug-likeliness and synthesizability. Our goal is to learn a molecule generation policy that constructs molecules $L$ given protein structure $P$ with probability matching their reward (exponentiated with $\beta$), such that $\pi(L|P,\beta)\propto R(L|P)^{\beta}$.
Method overview. The protein pocket is first featurized as a $K$-nearest neighbor residue graph and encoded using GVP-GNN (Jing et al., 2021) (section: 4.1). Then, we use this pocket embedding as the condition for GFlowNet molecule generation (section 4.2). Finally, we reward generated molecules using our docking score prediction model which leverages protein-ligand interaction priors (section 4.3). The high-level architecture of our SBDD framework is illustrated in Figure 1.
@@ 4.2 Pocket conditioned GFlowNet
Molecule representation. During molecular generation, we represent ligands as a 2D molecular graph $\mathcal{G}^{L}=(\mathcal{V}^{L},\mathcal{E}^{L})$ with node $v^{L}_{i}\in\mathcal{V}^{L}$ representing a molecule fragment, and directional edges $e^{L}_{ij}\in\mathcal{E}^{L}$ indicating the attachment atom of fragment $v^{L}_{i}$ that connects to fragment $v^{L}_{j}$. Since a molecule’s reward (representing its desirability as a real-world drug) should be the same regardless of its predicted 3D conformation, we represent ligands as 2D graphs here.  [^1] Docking score is dependent on a molecule’s 3D conformation. However, we consider the best docking score for a molecule as the reward here, which is invariant to its conformation in this context. Taking the best docking score is a valid approach, because when a compound is tested for binding in the wet lab, the compound simply binds the protein with the conformations which results in the strongest affinity. Furthermore, the predicted 3D position of generated molecules from existing diffusion-based SBDD models often changes significantly upon re-docking and likely do not reflect the true conformation. (Reidenbach, 2024).
Fragment vocabulary construction. TacoGFN generates molecules by adding one molecular fragment at a time. To create the vocabulary of fragments used, we extract common fragments from a chemical database in a data-driven and chemically valid way. To obtain a fragment vocabulary, we first apply BRICS decomposition (Degen et al., 2008) to 250k ZINC20 (Irwin et al., 2020) molecules. BRICS breaks molecules via retrosynthetic rules and provides synthetically accessible building blocks, i.e. molecules that are easy to prepare (Seo et al., 2023). To reduce the fragment vocabulary size, we further break all single bonds connecting a heavy atom to a ring structure. Next, we retain the fragments that occur in more than 50 (or 0.02%) of the molecules in our ZINC set. Only atoms (of a fragment) in the bonds decomposed by the BRICS can join to form new fragment connections during generation time; This reduces the occurrence of non-synthesizable attachment points forming bonds. This results in a set of 475 building block-like fragments. As our fragments are mined from a synthetically accessible virtual library and poor synthetic accessibility is penalized in the reward function, TacoGFN achieves superior synthetic accessibility compared to all existing SBDD methods.
Molecular generation framework. We formulate molecular generation as a sequential decision process and implement it using a GFlowNet. At the $t$-th step, the forward action policy $P_{a}$ samples an action $a$ depending on the current molecule state $\mathcal{G}^{L}_{t}$ and pocket embedding $\mathbf{h}_{\mathcal{G}^{\mathcal{P}}}$. The transition function $\mathcal{T}$ is a deterministic function which applies the action to the molecular graph at step $t$ to produce a molecule graph at step $t+1$.
$a_{t}$ $\sim P_{a}(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ (2)
@@
Previous autoregressive models often formulate action prediction as a supervised task, where the goal is to predict the correct ground truth actions obtained from masked molecules (Peng et al., 2022). Instead, our molecule generation policy $P_{a}$ aims to generate molecules with probability $P(\mathcal{G}^{L}|\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ proportional to the reward.
We use a graph transformer (Yun et al., 2020) to model the probability $P_{a}(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$, by taking the partially constructed molecular graph at the $t$-th time step $\mathcal{G}^{L}_{t}$ and the pocket embedding $\mathbf{h}_{\mathcal{G}^{\mathcal{P}}}$ as input. The input feature $\mathbf{h}^{L(0)}_{i}$ of molecular node $v^{L}_{i}$ is a one-hot encoding of the node’s fragment type. The molecular edge input feature $\mathbf{e}_{ij}^{L(0)}$ is a one-hot mapping of the attachment atom index in node $v^{L}_{i}$ which connects to node $v^{L}_{j}$. We add an additional virtual conditioning node $\mathbf{h}^{V(0)}$, featurized using pocket embedding $\mathbf{h}_{\mathcal{G}^{\mathcal{P}}}$, to the graph $\mathcal{G}^{L}_{t}$ (Pham et al., 2017). This virtual node is connected to all other nodes and serves as a graph-level node to provide pocket information. After $N$ Transformer layers, the set of final node embeddings $\{\mathbf{h}^{L(N)}_{i}\}$ and edge embeddings $\{\mathbf{e}^{L(N)}_{ij}\}$ are obtained. The final graph level embedding $\mathbf{g}^{L(N)}$ is obtained via the concatenation of the global average pooling of the node embeddings and the final virtual node embedding $\mathbf{h}^{V(N)}$, as seen in Equation 4.
Using these final molecular graph embeddings, we follow the actions defined in previous works on fragment-based molecular generation (Bengio et al., 2021; Jin et al., 2018). Full details can be found in Appendix C.1.1.
Reward function. A drug candidate must not only have a high affinity to the pocket but also satisfy drug-like properties and synthetizability requirements to be selected for experimental validation. We design a reward function by multiplying the normalized score from all three aspects, using QED (Bickerton et al., 2012) as a measure of drug-likeliness, SA (Ertl & Schuffenhauer, 2009) as a measure of ease-of-synthesizability, and predicted docking score as a measure of affinity between ligand and protein [^2] Computing the reward is very fast, for example, computing it for a batch of 64 protein-ligand pairs took under 0.15 seconds.. In drug discovery, it is crucial to design molecules that simultaneously satisfy all of these relevant properties, despite the inherent trade-offs among these properties. By multiplying the scores, the reward function ensures that a low score in any one of the factors (QED, SA, or DS) will significantly reduce the overall reward. Therefore, multiplying the scores is a more appropriate choice than summing the scores here.
We just have to optimize QED and SA up to a certain threshold for a molecule to be suitable as a drug candidate - optimizing them further does not bring additional utility (Coley, 2021). Therefore, we clip the reward for the QED or SA component to 1 when they achieve their respective threshold $t_{QED}$ and $t_{SA}$. For example, the reward function will prioritize more on optimizing QED if $t_{QED}$ is higher, at the expense of other properties such as affinity (See details of reward function in Appendix C.1.2).